OpenStudy (anonymous):

how does csc(-120) equal to -csc(120)?

Nnesha (nnesha):

sin and tan are odd cos = even \[\large\rm sin(-\theta)= -\sin \theta\]\[\large\rm cos(-\theta)= \cos \theta\]\[\large\rm Tan(-\theta)= - Tan \theta\]

OpenStudy (anonymous):

I still don't understand

Nnesha (nnesha):

what is reciprocal of sin ,cos and tan ???

OpenStudy (anonymous):

Honestly I have no idea. I am the worst with trigonometry. I don't understand it whatsoever

Nnesha (nnesha):

then u need to start it over \[\large\rm sin =\frac{ 1 }{ \csc }~~\cos=\frac{ 1 }{ \sec }~~~\tan=\frac{ 1 }{ \cot }=(\frac{ \sin }{ \cos })\] reciprocal of sin is 1/csc reciprocal of cos is 1/sec reciprocal of tan is 1/cost or you can say sin /cos csc=1/sin sec=1/cos cot=1/tan csc , sec and cot are reciprocal of sin , cos and tan

Nnesha (nnesha):

if sin =a/b then csc = b/a

Nnesha (nnesha):

so if sin is odd function then csc is odd or even ?

OpenStudy (anonymous):

odd

Nnesha (nnesha):

yes right that's how \[\huge\rm csc(-120)= - \csc(120)\]

OpenStudy (dumbcow):

another way of looking at it is using the unit circle and fact that neg angle goes clock-wise 

OpenStudy (anonymous):

but I'm just not understanding how the negative comes out

OpenStudy (dumbcow):

sin refers to "y-value" , on the circle sin(-120) is the neg y-value of sin(120)

OpenStudy (dumbcow):